# Anomalous and nonlinear Hall effects in two-dimensional CrPS4

### v1

@@ VI Layer Hall effect in bilayer
Figure 7: (a) Calculated layer Hall conductivity in bilayer CrPS4. Here, we focus on the peak structure in valence band. (b) Schematic figure showing the application of a gate electric field to bilayer CrPS4. This breaks the $\mathcal{PT}$ symmetry and generates a finite anomalous Hall response. (c) Layer-resolved Hall conductivities under a gate field of $0.02$ V/nm. (d) The resulting anomalous Hall conductivity versus gate field at a fixed chemical potential of $-0.25$ eV.
@@
In Fig. 7(a), we show the calculated layer Hall conductivity for bilayer CrPS4. One can see its peak value can reach $\sim 220$ S/cm on the hole doping side. This indicates that although the net charge current flow vanishes, the AHE response within each layer is still quite large. Therefore, if one can break the symmetry that connects the two layers (here it is $\mathcal{PT}$), it will be possible to generate an imbalance between $\sigma_{xy}^{T}$ and $\sigma_{xy}^{B}$ and induce a finite AHE. A simplest method is by using the gate electric field.
Figure 7(c) shows the calculated layered-resolved and net AHE conductivities for bilayer CrPS4 under a gate field of $0.02$ V/nm. One observes that the gate field shifts the layered-resolved conductivities $\sigma_{xy}^{\ell}$ in different ways, leading to a finite AHE. In Fig. 7(d), we plot the net AHE conductivity as a function of gate field at a fix chemical potential $\mu=-0.25$ eV. One observes the linear scaling at weak field region, and the sign of AHE response can be controlled by the direction of gate field, which is a nice feature for device applications.
@@ VII In-plane anomalous Hall effect in bilayer
Another way to break the $\mathcal{PT}$ symmetry of bilayer CrPS4 is to apply an in-plane magnetic field [55]. (Out-of-plane $B$ field also works, but it will bring in the ordinary Hall effect.) The main effect of this in-plane $B$ field is to tilt the local magnet moments and cause a net in-plane magnetization $\delta M$ along the $B$ field direction. It was found from experiment that the magnetic anisotropy in CrPS4 is relatively weak, so the spin canting can be readily achieved by moderate $B$ field strength [29].
@@
The direction of in-plane $B$ field affects the induced AHE. For field along $y$ direction, the resulting magnetization $\delta M_{y}$ cannot efficiently generate AHE, because it preserves the $\mathcal{M}_{y}$ symmetry for each layer. In comparison, the component $\delta M_{x}$ along $x$ direction is more effective in producing the AHE response. In the following, we shall focus on the this case, with $B$ field causing spin canting in the $x$ direction.
In DFT calculation, we model the spin-canted state by rotating the local moments towards the $x$ direction by a canting angle $\theta$. At a small angle $\theta=30^{\circ}$, the calculated layered-resolved and net AHE conductivities are plotted in Fig. 8(c). One observes that the contributions from the two layers no longer cancel out. A sizable AHE conductivity with peak value $\sim 132$ S/cm can be obtained. In Fig. 8(b), we plot the induced AHE conductivity as a function of angle canting $\theta$ for a fixed $\mu=-0.31$ eV. The curve shows a linear increase at small angles and tends to saturate at large angle. It reaches $\sim 201$ S/cm at angle $\theta=60^{\circ}$.
The magnitude of $B$ field needed to achieve the canting angle $\theta$ can be estimated as $B\approx 2J_{c}S\sin(\theta)/(g\mu_{\mathrm{B}})$. Here, we take the interlayer exchange coupling $J_{c}=0.15$ meV as obtained from experiment [25], and $S=3/2$. In this estimation, the magnetic anisotropy energy is neglected, since it is two order of magnitude smaller than the exchange energy [25]. We find that a canting angle of $5^{\circ}$ can be achieved by a moderate $B$ field of $340$ mT, showing that using in-plane $B$ field is an effective way to control the AHE response in bilayer CrPS4.
@@
Figure 8: (a) Schematic figure showing the application of an in-plane magnetic field to bilayer CrPS4. The field breaks $\mathcal{PT}$ symmetry, tilts the local moments, and induces a finite linear anomalous Hall response. (b) Calculated anomalous Hall conductivity versus spin-canting angle $\theta$ at a fixed chemical potential of $-0.31$ eV. (c) Calculated layer-resolved and total anomalous Hall conductivities for the bilayer when the spin-canting angle $\theta=30^{\circ}$.
@@ VIII Discussion and Conclusion
In this study, we have investigated the rich anomalous Hall transport phenomena in 2D CrPS4 systems. Monolayer and bilayer CrPS4 are used as representatives for demonstration. The general features apply also to other few-layer CrPS4 systems. The odd layers share the same symmetry as monolayer, while the even layers share the same symmetry as bilayer. Hence, linear AHE is expected in odd layers, whereas nonlinear Hall effect is expected in even layers. The linear AHE induced by gate field or in-plane magnetic field should occur in even-layer CrPS4 as well.
@@
Experimentally, 2D CrPS4 has been successfully fabricated via various techniques down to monolayer limit [25, 28]. Field effect transistor devices based on 2D CrPS4 have also been demonstrated [37, 31, 36, 35]. The phenomena predicted here should be detectable via standard transport measurement with a dual gate device setup. For the nonlinear Hall effect, the typical method is to modulate the driving current with a low frequency (less than 100 Hz) and measure the response signal at second harmonic frequency by using the lock-in technique [11, 12]. The angular dependence of the signal (as manifested in Eq. (9)) can be probed by fabricating a disk shaped device with multiple leads to allow driving current (and response signal) exerted (detected) along different in-plane directions.
@@
For odd-layer CrPS4, the nonlinear transport signal should vanish both above and below magnetic transition temperature $T_{c}$. For even layers, the nonlinear signal is forbidden above $T_{c}$ (since the crystal structure preserves $\mathcal{P}$) and appears only when temperature is lowered below $T_{c}$. This could be used as a signature for detecting $T_{c}$ or the parity of number of layers. In addition, we note that for temperature above $T_{c}$, both linear and second-order Hall effects are forbidden by the $\mathcal{T}$ and $\mathcal{P}$ symmetries. The leading order anomalous Hall response should be from the third order [51, 56], which was recently probed in several nonmagnetic materials.
@@
In this work, we focus on the intrinsic responses, which represent inherent material properties. There also exist extrinsic contributions arising from disorder scattering, which vary from sample to sample [57]. Experimentally, a typical method to distinguish the different contributions is to perform a scaling analysis by varying the various disorder strengths. The intrinsic response correspond to a component that is independent of scattering. And first-principles result serves as an important benchmark for such analysis.
@@
In conclusion, we have investigated various anomalous Hall transport phenomena in 2D CrPS4. Due to their low crystal symmetry and layer-dependent magnetism, we find interesting linear and nonlinear AHEs, with sizable conductivities predicted by our first-principles calculations. These effects allow us to probe the intrinsic band geometric quantities, such as Berry curvature, BCP, and quantum metric, of the material system. The demonstrated controllability of Hall transport in 2D CrPS4 also offers possible mechanisms for device operations. Our predications here can be readily tested in experiment, and they may find promising applications in next-generation electronics and spintronics based on 2D magnetic materials.
@@ Acknowledgements.
The authors thank D. L. Deng for valuable discussions. This work is supported by UM Start-up Grant (SRG2023-00033-IAPME, SRG2022-00030-IAPME, and SRG2023-00057-IAPME), UM Multi-Year Research Grant (MYRG-GRG2023-00206-IAPME-UMDF), Science and Technology Development Fund of Macau SAR (0048/2023/RIB2 and 0066/2024/RIA1), National Key R&\&&D Program of China (Grant No. 2022YFA1402600), and NSFC (Grants No. 12174018).

### v2

@@ VI Layer Hall effect in bilayer
Figure 7: (a) Calculated layer Hall conductivity in bilayer CrPS4. Here, we focus on the peak structure in valence band. (b) Schematic figure showing the application of a gate electric field to bilayer CrPS4. This breaks the $\mathcal{PT}$ symmetry and generates a finite anomalous Hall response. (c) Layer-resolved Hall conductivities and total Hall conductivities under a gate field of $0.02$ V/nm. (d) The resulting anomalous Hall conductivity versus gate field at a fixed chemical potential of 0.25-0.25 eV.
In Fig. 7(a), we show the calculated layer Hall conductivity for bilayer CrPS4. One can see its peak value can reach $\sim 220$ S/cm on the hole doping side. This indicates that although the net charge current flow vanishes, the AHE response within each layer is still quite large, leading to a large layer Hall effect.
It is noted that the existence of layer Hall effect does not require the breaking of $\mathcal{PT}$ symmetry. Nevertheless, direct detection of layer resolved Hall response is difficult. To probe layer Hall effect in experiment, an indirect way is to break the $\mathcal{PT}$ symmetry that connects the two layers by gating, so that the layer Hall effect can be converted to an anomalous Hall signal [76].Figure 7(c) shows the calculated layered-resolved and net AHE conductivities for bilayer CrPS4 under a gate field of $0.02$ V/nm (a moderate gate field strength that can be readily achieved in experiment). One observes that the gate field shifts the layered-resolved conductivities $\sigma_{xy}^{\ell}$ in different ways, leading to a finite AHE. In Fig. 7(d), we plot the net AHE conductivity as a function of gate field at a fix chemical potential $\mu=-0.25$ eV. One observes the linear scaling at weak field region, and the sign of AHE response can be controlled by the direction of gate field, which is a nice feature for device applications.
@@ VII In-plane anomalous Hall effect in bilayer
As mentioned above, even-layer CrPS4 should have vanishing linear AHE due to the $\mathcal{PT}$ symmetry. By applying an in-plane magnetic field, the $\mathcal{PT}$ symmetry can be broken, and a linear charge Hall signal can be produced [78, 79, 80, 81]. (Out-of-plane $B$ field also works, but it will bring in the ordinary Hall effect.) In literature, such a configuration is commonly called the planar Hall effect. However, it should be noted that in many previously reported experiments, the transverse signal is even in the applied magnetic field $B$, which is actually not a Hall response but a transverse magneto-resistance. This is in contrast to the effect considered here, which is odd in $B$. Because of this difference, some recent experimental studies termed this effect as in-plane AHE [82, 83, 84], which we adopt here.
@@
The main effect of applied in-plane $B$ field here is to tilt the local magnet moments and cause a net in-plane magnetization $\delta M$ along the $B$ field direction. This picture was confirmed in experiment, and it was found that the magnetic anisotropy in CrPS4 is relatively weak, so the spin canting can be readily achieved by moderate $B$ field strength [30]. The direction of in-plane $B$ field affects the induced AHE. For field along $y$ direction, the resulting magnetization $\delta M_{y}$ cannot efficiently generate AHE, because it preserves the $\mathcal{M}_{y}$ symmetry for each layer. In comparison, the component $\delta M_{x}$ along $x$ direction is more effective in producing the in-plane AHE response. In the following, we shall focus on this case, with $B$ field causing spin canting in the $x$ direction.
@@
Figure 8: (a) Schematic figure showing the application of an in-plane magnetic field to bilayer CrPS4. The field breaks $\mathcal{PT}$ symmetry, tilts the local moments, and induces a finite linear anomalous Hall response. (b) Calculated anomalous Hall conductivity versus spin-canting angle $\theta$ at a fixed chemical potential of 0.31-0.31 eV. (c) Calculated layer-resolved and total anomalous Hall conductivities for the bilayer when the spin-canting angle $\theta=30^{\circ}$.
In DFT calculation, we model the spin-canted state by rotating the local moments towards the $x$ direction by a canting angle $\theta$. At a small angle $\theta=30^{\circ}$, the calculated layered-resolved and net AHE conductivities are plotted in Fig. 8(c). One observes that the contributions from the two layers no longer cancel out. A sizable AHE conductivity with peak value $\sim 132$ S/cm can be obtained. In Fig. 8(b), we plot the induced AHE conductivity as a function of canting angle $\theta$ for a fixed $\mu=-0.31$ eV. The curve shows a linear increase at small angles and tends to saturate at large angle. It reaches $\sim 201$ S/cm at angle $\theta=60^{\circ}$. Using a simple AFM model, the magnetic energy under an in-plane magnetic field can be written as $E_{m}=-2g\mu_{B}BS\sin\theta-J_{c}S^{2}\cos 2\theta$, where $S$ is the magnitude of local spin, and $J_{c}$ is the interlayer AFM coupling strength [85]. In this estimation, the magnetic anisotropy energy is neglected, since it is two order of magnitude smaller than the exchange energy [25]. Minimizing this energy with respect to $\theta$, we find the magnitude of $B$ field needed to achieve canting angle $\theta$ can be estimated as
$B\approx 2J_{c}S\sin(\theta)/(g\mu_{\mathrm{B}}).$ (15)
Taking the interlayer exchange coupling $J_{c}=0.15$ meV as obtained from experiment [25], and $S=3/2$, we find that a canting angle of $5^{\circ}$ can be achieved by a moderate $B$ field of $340$ mT, showing a pronounced in-plane AHE response in bilayer CrPS4.
@@ VIII Discussion and Conclusion
In this study, we have investigated the rich anomalous Hall transport phenomena in 2D CrPS4 systems. Monolayer and bilayer CrPS4 are used as representatives for demonstration. The general features apply also to other few-layer CrPS4 systems. The odd layers share the same symmetry as monolayer, while the even layers share the same symmetry as bilayer. Hence, linear AHE is expected in odd layers, whereas intrinsic nonlinear Hall effect is expected in even layers. The linear AHE induced by gate field or in-plane magnetic field should occur in even-layer CrPS4 as well. It is worth noting that the nonlinear Hall response would vanish in the bulk limit, since bulk CrPS4 retains inversion symmetry regardless of the direction of its Néel vector.
@@
Experimentally, 2D CrPS4 has been successfully fabricated via various techniques down to monolayer limit [25, 28]. Field effect transistor devices based on 2D CrPS4 have also been demonstrated [35, 39, 40, 44]. The phenomena predicted here should be detectable via standard transport measurement with a dual gate device setup. For the nonlinear Hall effect, the typical method is to modulate the driving current with a low frequency (less than 100 Hz) and measure the response signal at second harmonic frequency by using the lock-in technique [11, 12]. The angular dependence of the signal (as manifested in Eq. (9)) can be probed by fabricating a disk shaped device with multiple leads to allow driving current (and response signal) exerted (detected) along different in-plane directions. As for the layer Hall effect in bilayer, besides detection via charge Hall signal under electric gating, it induces an in-plane orbital magnetization [77] and hence may also be probed by magneto-optical Kerr microscopy [19, 22].
@@
For odd-layer CrPS4, the nonlinear transport signal should vanish both above and below magnetic transition temperature $T_{c}$. For even layers, the nonlinear signal is forbidden above $T_{c}$ (since the crystal structure preserves $\mathcal{P}$) and appears only when temperature is lowered below $T_{c}$. This could be used as a signature for detecting $T_{c}$ or the parity of number of layers. In addition, we note that for temperature above $T_{c}$, both linear and second-order Hall effects are forbidden by the $\mathcal{T}$ and $\mathcal{P}$ symmetries. The leading order anomalous Hall response should be from the third order  [7, 74, 86, 87, 88, 89, 90, 91], which was recently probed in several nonmagnetic materials.
@@
In this work, we focus on the intrinsic responses, which represent inherent material properties. There also exist extrinsic contributions arising from disorder scattering, which vary from sample to sample [92, 93, 72]. Experimentally, a typical method to distinguish the different contributions is to perform a scaling analysis by varying the various disorder strengths. The intrinsic response correspond to a component that is independent of scattering. And first-principles result serves as an important benchmark for such analysis. As we have mentioned, even-layer CrPS4 represents a good platform for studying intrinsic nonlinear AHE, because of the following reasons: (1) It low-symmetry monoclinic structure makes the intrinsic nonlinear AHE a property of the magmatic material itself, without requiring interfacial effects as in previous experiments; (2) Its spacetime inversion symmetry suppresses the extrinsic contributions, making the intrinsic response more important; and (3) its 2D vdW character makes the system highly tunable and readily integrated into vdW heterostructures for nonlinear device applications.
@@
In conclusion, we have investigated various anomalous Hall transport phenomena in 2D CrPS4. Due to their low crystal symmetry and layer-dependent magnetism, we find a rich variety of anomalous Hall responses, with sizable conductivities predicted by our first-principles calculations. These effects allow us to probe the intrinsic band geometric quantities, such as Berry curvature, BCP, and quantum metric, of the material system. The demonstrated controllability of Hall transport in 2D CrPS4 also offers possible mechanisms for device operations. Our predications here can be readily tested in experiment, and they may find promising applications in next-generation electronics and spintronics based on 2D magnetic materials.
@@ Acknowledgements.
The authors thank D. L. Deng for valuable discussions. C.X. acknowledges the support by the start-up funding from Fudan University. This work is supported by UM Start-up Grant (SRG2023-00033-IAPME, SRG2022-00030-IAPME, and SRG2023-00057-IAPME), UM Multi-Year Research Grant (MYRG-GRG2023-00206-IAPME-UMDF), Science and Technology Development Fund of Macau SAR (0048/2023/RIB2 and 0066/2024/RIA1), National Key R&\&D Program of China (Grant No. 2022YFA1402600), NSFC (Grants No. 12174018) and HK PolyU Start-up Grant (P0057929).